# Riemann-Silberstein geometric phase for high-dimensional light manipulation

### v1

@@ Results
The output wavefronts can be further increased by setting $\alpha_{\mathrm{p}}=\alpha_{\mathrm{m}}\neq\alpha_{\mathrm{rs}}$ for the RS metasurface in Fig. 4a. Under the incidence of the plane waves $\left|\sigma_{\mathrm{in}}=\pm 1,\kappa_{\mathrm{in}}=\pm 1\right\rangle$, this metasurface can generate twelve distinct output wavefronts propagating in different directions, as schematically shown in Fig. 5a. The simulated far-field intensity patterns are shown in Fig. 5b, where the intensity lobes are labelled in accordance with the numbers in Fig. 5a. As seen, the scattered far fields can propagate in twelve different directions, depending on the values of $\sigma_{\text{out}}$ and $\kappa_{\text{out}}$. Additionally, the number of distinct outputs can be dynamically reconfigured by employing the incident wave $\left|\Psi_{\text{in}}\right\rangle=\eta_{1}|+1,+1\rangle+\eta_{2}|+1,-1\rangle+\eta_{3}|-1,+1\rangle+\eta_{4}|-1,-1\rangle$ with varying 4D polarization. For example, by tuning the coefficients $\eta_{1-4}$ to obtain four different incident polarizations, denoted by the points “i-iv” on the Poincaré hypersphere in Fig. 5c, the metasurface can generate 12, 10, 8, and 6 far-field intensity lobes, as shown in Fig. 5d. These results demonstrate that the general RS metasurface can enable high-dimensional wavefront manipulations beyond the capabilities of conventional metasurfaces. The functionalities can be further enriched if time-varying or active components are introduced into the system [58, 59, 60, 61].
@@ Discussion
In summary, we unravel a high-dimensional manifestation of the geometric phase induced by complete electromagnetic polarization evolution in the 4D RS Hilbert space. Beyond the conventional PB phase induced by electric or magnetic polarization evolution, we discover a new geometric phase—RS phase, which originates from the hybrid electric-magnetic polarization evolution and can emerge even in linearly polarized waves. The findings provide fundamental insights into the geometric nature of high-dimensional light-matter interactions and enable multiplexed wavefront shaping through the RS metasurfaces. The proposed mechanism applies to general electromagnetic waves including evanescent waves and complex structured waves. Our work broadens the geometric phase paradigm and introduces new degrees of freedom for light manipulation, which can facilitate multifunctional nanophotonic design and enhance communication capacity.
@@ Acknowledgements
The work described in this paper was supported by grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (Projects Nos. AoE/P-502/20 and CityU11308223) and National Natural Science Foundation of China (No. 12322416). D.P.T. acknowledges support from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Nos. C5031-22G, C5078-24G, CityU11305223, CityU11300224, CityU11304925, and CityU11305125), City University of Hong Kong (Project No. 9380131), and National Natural Science Foundation of China (Grant No. 62375232). G.-B.W. acknowledges support from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. CityU21207824). The authors thank Profs. C. T. Chan and Z. Q. Zhang for helpful discussions. The authors also thank Dr. Ka Fai Chan and Dr. Chenfeng Yang for their support in the experiments.

### v2

@@ Introduction
The output wavefronts can be further increased in the higher-order RS metasurface with $\alpha_{\mathrm{p}}=\alpha_{\mathrm{m}}\neq\alpha_{\mathrm{rs}}$. Under the incidence of the plane waves $\left|\sigma_{\mathrm{in}}=\pm 1,\kappa_{\mathrm{in}}=\pm 1\right\rangle$, this metasurface can generate twelve output wavefronts propagating in different directions, as schematically shown in Fig. 5a. The simulated far-field intensity patterns are shown in Fig. 5b, where the intensity lobes are labelled in accordance with the numbers in Fig. 5a. As seen, the scattered far fields can propagate in twelve different directions, depending on the values of $\sigma_{\text{out}}$ and $\kappa_{\text{out}}$. Additionally, the number of distinct outputs can be dynamically reconfigured by employing the incident wave $\left|\Psi_{\text{in}}\right\rangle=\eta_{1}|+1,+1\rangle+\eta_{2}|+1,-1\rangle+\eta_{3}|-1,+1\rangle+\eta_{4}|-1,-1\rangle$ with varying 4D polarization. For example, by tuning the coefficients $\eta_{1-4}$ to obtain four different incident polarizations, denoted by the points “i-iv” on the Poincaré hypersphere in Fig. 5c, the metasurface can generate 12, 10, 8, and 6 far-field intensity lobes, as shown in Fig. 5d. These results demonstrate the practical application potential of the framework, which can be further enriched by introducing time-varying or active components into the system [zhang2018space, shaltout2019spatiotemporal, galiffi2022photonics, tirole2023double].
@@ Discussion
We unveil the geometric phases arising from complete electromagnetic polarization evolution in the 4D RS space by treating electromagnetic waves as a four-component bispinor. Beyond the conventional PB phase induced by electric or magnetic polarization evolution, we discover a new geometric phase, the RS phase, which originates from the hybrid electric-magnetic polarization evolution and can manifest even in linearly polarized waves. The complementary RS and PB phases enable a unified high-dimensional geometric framework for understanding and controlling phase shifts in light propagation across general interfaces, including artificial metasurfaces. The proposed mechanism applies to general electromagnetic waves, including evanescent waves and complex structured waves. Our work broadens the geometric phase paradigm and offers fundamental insights into the geometric nature of light-matter interactions, opening new avenues for exploring topological and non-Abelian phenomena in high-dimensional classical wave systems. Future research may integrate the spin-redirection geometric phase [chiao1986manifestations, RN45] arising from variations of propagation direction into this framework, which promises a more comprehensive approach for geometric phase physics.
After the initial posting of this work [cheng2025riemann], a related preprint appeared [vernon2025electric], proposing an “electric-magnetic (EM) geometric phase” arising from cyclic evolution of the $\mathbf{E}$$\mathbf{H}$ relationship in nonparaxial light. This EM geometric phase is conceptually equivalent to the RS geometric phase described here and can be regarded as a special case within our broader high-dimensional framework based on the 4D RS space. The RS geometric phase applies to general electromagnetic waves, including nonparaxial waves (e.g., evanescent waves and structured waves). Their conclusion that EM geometric phase appears exclusively in nonparaxial light does not conflict with our results, as their analysis does not address the scenarios involving material interfaces, where the local $\mathbf{E}$$\mathbf{H}$ relationship can be modified. We view their work as complementary to ours, together contributing to a more complete understanding of this geometric phase across different wave regimes.
@@ Acknowledgements
The work described in this paper was supported by grants from the Research Grants Council of the Hong Kong Special Administrative Region, China (Projects Nos. AoE/P-502/20 and CityU11308223) and National Natural Science Foundation of China (No. 12322416). D.P.T. acknowledges support from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project Nos. C5031-22G, C5078-24G, CityU11305223, CityU11300224, CityU11304925, and CityU11305125), City University of Hong Kong (Project No. 9380131), and National Natural Science Foundation of China (Grant No. 62375232). G.-B.W. acknowledges support from the Research Grants Council of the Hong Kong Special Administrative Region, China (Project No. CityU21207824). The authors thank Prof. C. T. Chan and Prof. Z. Q. Zhang for helpful discussions. The authors also thank Dr. Ka Fai Chan and Dr. Chenfeng Yang for their support in the experiments.
Author contributions
S.W. and Y.C. conceived the idea and developed the concepts. Y.C. designed the structures and conducted the numerical simulations. Y.-S.Z., Y.C., and G.-B.W. designed and carried out the experiments. Y.C. and S.W. wrote the draft. S.W., G.-B.W., and D.P.T. supervised the project. All authors contributed to discussions, interpretation of the results, and polishing of the manuscript.